OpenStudy (anonymous):

can someone help me understand this step, i am finding the second derivatve using implicit differ. : (y-(x)(-x/y))/(y^2), how do i simpify this?

OpenStudy (anonymous):

\[\frac{y-x\times \frac{-x}{y}}{y^2}\] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i am finding the second derivative and i got this far, but i am unsure of how to simplfy in order to get it in the form needed

OpenStudy (anonymous):

multiple x times (-x/y) and get x^2/y, then find a common denominator inorder to add the two fractions in the numerator. Finally multply by the reciprocal or y^2 which is 1/y.

OpenStudy (anonymous):

1/y^2*

OpenStudy (amistre64):

my next goal should be to randomly colorize equations :)